OpenStudy (anonymous):

sinx/(sinx+cosx)=tanx/(1+tanx)

OpenStudy (anonymous):

\[\frac{ sinx}{ sinx+cosx} = \frac{ tanx }{ 1+tanx }\]

OpenStudy (anonymous):

\[\frac{ sinx + \cos x }{\sin x } = \frac{1 + tanx}{tanx}\] take the inverse of each side

OpenStudy (anonymous):

hint: \[1+ \tan x = \frac{ \sin x + \cos x }{\cos x} \] as \[\tan x = \frac{ \sin x }{ \cos x }\] try it yourself/